What number when multiplied by itself is 13 greater than the preceding number when it is multiplied by itself

Answer by Aldorozos(172)

x*x = 13 +(x-1)(x-1) =
x^2 = 13 + (x^2-2x - 1)
x^2 =13 + x^2 -2x +1
x^2 cancels out from both sides
We end of having -2x +14 = 0
Then x = 7 and the preceding number is 6
We know that if we multiply 7*7 = 13 + 6*6